OpenStudy (anonymous):

if f(x) = x/3 and g(x) = 1/x find f(x) - g(x)

OpenStudy (anonymous):

\[ f(x) - g(x) =\frac x3- \frac 1x \]

OpenStudy (anonymous):

They probably want you to simplify it